1. A sample of gaseous uranium hexafluoride, UF6, is held at a temperature of 300 K and a pressure of 0.1 mbar. The collision diameter of UF6 is 0.40 nm. (i) Calculate the root-mean-square (r.m.s.) speed of the molecules. Make sure to show all units in your working and how the units cancel out. [5 marks] (ii) Estimate the mean free path and collision frequency under these conditions. [10 marks]


1
Expert's answer
2022-02-20T08:36:17-0500

"V_{rms}=\sqrt{\frac{3RT}{M}}"


Molar mass of "UF_6=352g\/mol"

"\therefore\>M=0.
352kg/mol"




"R=8.314J\/mol \/K"

"T=300K"



"V_{rms}=\frac{\sqrt{3×8.314Jmol^{-1}K^{-1}300K}}{0.352kgmol^{-1}}"



"=145.8\sqrt{Jkg^{-1}}"


But "J=kgm^2s^{-2}"

"=145.8\sqrt{kgm^2s^{-2}kg^{-1}}"


"=145.8m\/s"


Part (ii)


Mean free path "\lambda" is given by


"\lambda=\frac{KT}{\sqrt{2}\pi\>d^2P}"


"=\frac{1.38×10^{-23}×300}{\sqrt{2}×\pi×(0.4×10^{-9})^2×0.1×10^2}"


"=5.824\u00d710^{-4}m"


Collision frequency


"=\frac{V_{rms}}{\lambda}=\frac{145.8}{5.824×10^{-4}}=2.503×10^5s^{-1}"